OpenStudy (anonymous):

NEED HELP FAST PLEASE!!! Problem: If you take an imagined number and add a zero to the right and subtract the result from 143, you'll get the tripled number. What is the imagined number.

OpenStudy (bahrom7893):

143-10x=3x 143 = 13x x = 143/13

OpenStudy (bahrom7893):

x=11

OpenStudy (anonymous):

....how did you do that? O.o

OpenStudy (bahrom7893):

was that fast enough?

OpenStudy (anonymous):

:P smartypants

OpenStudy (bahrom7893):

11, add a 0 on the right: 110 Subtract the result from 143: 143-110 = 33 and u'll get triple the number. 33 = 11*3, that's right..

OpenStudy (anonymous):

Thanks